For the arithmetic sequence with a1=3 and a20=117, find the common difference D.
Answer by Theo(13342)

the formula is:

An = A1 + (n-1) * d

An = 117
A1 = 3
n = 20

the formula becomes:

117 = 3 + 19 * d
subtract 3 from both sides of the equation to get:
114 = 19 * d
divide both sides of the equation by 19 to get:
114 / 19 = d
simplify to get:
6 = d which is the same as:
d = 6

your formula becomes:

117 = 3 + 19 * 6 which becomes:
117 = 3 + 114 which becomes:
117 = 117 which confirms your solution is correct.